# Towards Large Language Models with Self-Consistent Natural Language Explanations

### v2

@@ Abstract
Large language models (LLMs) seem to offer an easy path to interpretability: just ask them to explain their decisions. Yet, studies show that these post-hoc explanations often misrepresent the true decision process, as revealed by mismatches in feature importance. Despite growing evidence of this inconsistency, no systematic solutions have emerged—partly due to the high cost of estimating feature importance, which limits evaluations to small datasets. To address this, we introduce the Post-hoc Self-Consistency Bank (PSCB)—a large-scale benchmark of decisions spanning diverse tasks and models, each paired with LLM-generated explanations and corresponding feature importance scores. Analysis of PSCB reveals that self-consistency scores barely differ between correct and incorrect predictions. We also show that the standard metric fails to meaningfully distinguish between explanations. To overcome this limitation, we propose an alternative metric that more effectively captures variation in explanation quality. We use it to fine-tune LLMs via Direct Preference Optimization (DPO), leading to significantly better alignment between explanations and decision-relevant features—even under domain shift. Our findings point to a scalable path toward more trustworthy, self-consistent LLMs. [^1] Code and data will be released upon acceptance.
Towards Large Language Models with Self-Consistent Natural Language Explanations
Sahar Admoni1, Ofra Amir1, Assaf Hallak2, Yftah Ziser2 1Technion – Israel Institute of Technology 2Nvidia Research saharad@campus.technion.ac.il, oamir@technion.ac.il, {ahallak,yziser}@nvidia.com
@@ 1 Introduction
As Large language models (LLMs) are increasingly embedded in user-facing and decision-support systems, their outputs—and especially the explanations accompanying them—are often trusted by end users, even when independent verification of correctness or reasoning is impractical Sun et al. (2024). To foster trust, LLMs are often prompted to provide natural language explanations (NLEs) to articulate their reasoning Quan et al. (2024). However, these explanations frequently fail to reflect the model’s inner decision process, a shortcoming known as a lack of faithfulness Lyu et al. (2022). A related concept, self-consistency, measures the degree of alignment between the model’s decisions and the explanation, particularly in terms of feature importance. While self-consistency is not equivalent to faithfulness, it is widely viewed as a necessary prerequisite Parcalabescu and Frank (2023). Inconsistent explanations signal misalignment between reasoning and justifications, undermining trust in high-stakes domains like healthcare, law, or education. LLMs typically produce explanations through: (i) chain-of-thought reasoning, where the explanation is generated as part of the model’s reasoning process Wei et al. (2022), and (ii) justifications generated after decisions Wiegreffe et al. (2020) (Figure 1). We focus on the latter. Prior work has proposed methods to evaluate self-consistency in these settings Turpin et al. (2023); Lanham et al. (2023); Atanasova et al. (2023); Wiegreffe et al. (2020), but improving self-consistency remains an open challenge.
In this work, we take a step forward by introducing new benchmarks, metrics, and methods to enhance explanation consistency. We introduce the post-hoc self-consistency bank (PSCB) to systematically evaluate and improve explanation consistency across multiple LLMs and datasets. We compare two complementary consistency metrics and apply Direct Preference Optimization (DPO) Rafailov et al. (2023) to enhance consistency. Our contributions include:
  • Post-hoc Self-Consistency Bank (PSCB): We introduce PSCB, a large-scale benchmark for evaluating natural language explanation (NLE) self-consistency in LLMs, consisting of over 23,000 model decisions, each annotated with an attribution vector and paired with five explanations, each accompanied by its own attribution vector—totaling over 116,000 explanation instances—collected across four multiple-choice QA datasets and two instruction-tuned LLMs.
  • Empirical analysis of self-consistency: We conduct the first large-scale analysis of attribution-based self-consistency in LLMs and find that self-consistency is unaffected by whether the model’s answer is correct. We showcase the effectiveness of Spearman rank correlation as an alignment metric, highlighting its strong discriminative ability to capture quality differences across candidate explanations.
  • Improving self-consistency via preference supervision: We demonstrate that attribution-based preference data—automatically derived from our PSCB benchmark—can be used to fine-tune LLMs with Direct Preference Optimization (DPO), to improve self-consistency, both in-domain and with robust cross-domain generalization.
@@
Figure 2: Overview of the PSCB pipeline. Given a multiple-choice QA instance, the LLM generates an answer and associated attribution scores over the input. It is then prompted to produce $K$ diverse explanations (via temperature sampling) conditioned on the question and answer. Attributions are recorded for each explanation, and self-consistency scores are computed between decision and explanation attributions.
@@ 2 Background and Related Work
2.1 Feature Attribution Methods
Feature attribution methods estimate the contribution of input components (e.g., tokens) or internal elements (e.g., attention heads) and are central to interpreting model predictions Zhao et al. (2023). Attention-based methods are appealing due to their simplicity and low cost, but are widely criticized as unreliable indicators of importance (Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019). More robust alternatives include gradient-based methods, which use backpropagation to assign importance scores—e.g., Integrated Gradients Sundararajan et al. (2017) and LRP Bach et al. (2015); Montavon et al. (2019)—and perturbation-based methods, which modify inputs and observe output changes. Notable examples of the latter include input masking Petsiuk et al. (2018), Shapley value estimation via SHAP Lundberg and Lee (2017), and surrogate model fitting as used in LIME Ribeiro et al. (2016). While SHAP offers strong theoretical guarantees, it is computationally expensive, especially for large models. LIME, in contrast, provides a practical middle ground—offering informative explanations with much lower computational cost.
2.2 NLE Self-Consistency
A faithful natural language explanation (NLE) reflects the reasoning behind a model’s prediction, not just a plausible-sounding justification (Jacovi and Goldberg, 2020). However, LLMs often generate fluent explanations that diverge from their actual decision process (Narang et al., 2020; Turpin et al., 2023), prompting a shift toward evaluating self-consistency—the extent to which explanations rely on the same input features or reasoning patterns as the predictions. Self-consistency has emerged as a practical proxy for faithfulness, supported by diverse evaluation strategies. Perturbation-based methods test explanation robustness through counterfactual edits, noise injection, misleading feature bias, or corrupted chain-of-thought reasoning (Atanasova et al., 2023; Wiegreffe et al., 2020; Turpin et al., 2023; Lanham et al., 2023). Attribution-based approaches offer a more direct assessment by comparing feature importance vectors. For instance, CC-SHAP (Parcalabescu and Frank, 2023) evaluates consistency via Shapley value alignment between predictions and explanations. While precise and model-agnostic, such methods have been limited to small-scale datasets, raising questions about their scalability.
3 Post-hoc Self-Consistency Bank
This section introduces the Post-hoc Self-Consistency Bank. We describe the key components needed to construct each dataset: a base question-answering dataset, a target LLM, and a self-consistency metric. We then outline our design choices for building the full bank and conclude with key statistics and findings. Our pipeline (Figure 2) is inspired by CC-SHAP Parcalabescu and Frank (2023) and generalizes to diverse feature attribution methods and alignment metrics.
@@ 3.1 Sequence Feature Attribution
To analyze which input tokens influence a model’s generation behavior, we compute token-level attribution scores for sequence outputs. These scores identify the features most responsible for the model’s confidence in its predictions.
Input and Output Definition.
Let the input be a token sequence $\mathbf{x}=[x_{1},x_{2},\dots,x_{m}]$ of length $m$, and the output be a generated sequence $\mathbf{y}=[y_{1},y_{2},\dots,y_{n}]$ of length $n$. The model generates $\mathbf{y}$ autoregressively, producing a logit vector $\ell_{t}\in\mathbb{R}^{V}$ over the vocabulary $V$ at each decoding step $t$.
Output Scoring.
We define a scalar confidence score for the full sequence as the sum of log-probabilities: $\mathrm{score}(\mathbf{x},\mathbf{y})=\sum_{t=1}^{n}\log P(y_{t})$, where $\log P(y_{t})=\log\left(\frac{\exp(\ell_{t}[y_{t}])}{\sum_{v\in V}\exp(\ell_{t}[v])}\right).$ This scalar represents the target quantity for attribution: how much each token in $\mathbf{x}$ contributes to the overall model confidence in generating $\mathbf{y}$.
Attribution Computation.
We compute a token-level attribution vector over the input:
$\boldsymbol{\phi}^{(\mathbf{y})}=[\phi_{1}^{(\mathbf{y})},\phi_{2}^{(\mathbf{y})},\dots,\phi_{m}^{(\mathbf{y})}],$
where $\phi_{i}^{(\mathbf{y})}$ quantifies the contribution of token $x_{i}$ to the model’s confidence in generating $\mathbf{y}$. The attribution method—whether gradient-based, perturbation-based, or sampling-based—determines how $\boldsymbol{\phi}^{(\mathbf{y})}$ is computed. In our framework, this process is method-agnostic; specific techniques are described in Section 3.4.
Skip Token Filtering.
To avoid noise from irrelevant symbols, we define a set of skip tokens $\mathcal{S}\subseteq V$ (e.g., punctuation, special formatting). Tokens $x_{i}\in\mathcal{S}$ are excluded from the attribution process by holding them fixed during perturbation or masking steps. This focuses attribution on semantically meaningful inputs.
@@ 3.2 Measuring Self-Consistency
To evaluate whether a model’s explanation reflects the reasoning behind its decision, we assess the alignment between two attribution vectors derived from the same input $\mathbf{x}$:
  • $\mathbf{y}_{\mathrm{dec}}$: the model’s predicted answer or decision.
  • $\mathbf{y}_{\mathrm{exp}}$: a post-hoc natural language explanation generated by conditioning on both $\mathbf{x}$ and $\mathbf{y}_{\mathrm{dec}}$.
Let $\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})}$ and $\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}})}$ denote the attribution vectors over $\mathbf{x}$ for the decision and the explanation, respectively. We quantify the self-consistency using a similarity metric that measures alignment between the two attribution vectors $\mathcal{M}\left(\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})},\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}})}\right):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$. The choice of $\mathcal{M}$ determines the aspect of alignment being evaluated (e.g., magnitude, rank, overlap), and is discussed in detail in the following subsections.
@@ 3.3 Constructing Attribution-Based Preference Data
To support the evaluation and improvement of self-consistency in natural language explanations (NLEs), we construct a dataset where each instance contains both model outputs (decisions and explanations) and their corresponding input attribution vectors. This structure enables fine-grained, token-level comparisons between the model’s decision rationale and its post-hoc justifications.
Attribution-Enhanced Instances.
Each instance is represented as a structured quintuple:
$\left(\mathbf{x},\;\mathbf{y}_{\mathrm{dec}},\;\left\{\mathbf{y}_{\mathrm{exp}}^{(i)}\right\}_{i=1}^{k},\;\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})},\;\left\{\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}}^{(i)})}\right\}_{i=1}^{k}\right),$
where $\mathbf{x}$ is the original multiple-choice question, $\mathbf{y}_{\mathrm{dec}}$ is the model’s predicted answer to $\mathbf{x}$, $\left\{\mathbf{y}_{\mathrm{exp}}^{(i)}\right\}_{i=1}^{k}$ are $k$ post-hoc explanations conditioned on both $\mathbf{x}$ and $\mathbf{y}_{\mathrm{dec}}$, $\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})}$ is the attribution vector over $\mathbf{x}$ for the decision, and $\left\{\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}}^{(i)})}\right\}_{i=1}^{k}$ are attribution vectors for the explanations. This format allows us to compare explanation alignment across multiple candidate justifications.
Output Generation.
For each input $\mathbf{x}$, we prompt an instruction-tuned language model in a zero-shot setting to produce a decision $\mathbf{y}_{\mathrm{dec}}$. We then sample $k$ diverse post-hoc explanations $\mathbf{y}_{\mathrm{exp}}^{(i)}$ conditioned on both $\mathbf{x}$ and $\mathbf{y}_{\mathrm{dec}}$. This setup yields multiple plausible explanations for the same decision, enabling relative alignment comparison.
Attribution Computation.
We compute attribution vectors over the input $\mathbf{x}$ for both the decision and each explanation, based on the log-likelihood of the generated output sequence. Our framework is method-agnostic and supports a range of feature attribution techniques, including gradient-based and perturbation-based methods. Formatting and special tokens are excluded from attribution to focus on semantically meaningful content.
Alignment-Based Labeling.
We measure attribution alignment between each explanation and the decision using the metrics introduced in the previous subsection. Each explanation is scored based on its alignment with the decision, and we retain the highest- and lowest-scoring examples as $\mathbf{y}_{\mathrm{exp}}^{\text{chosen}}$ and $\mathbf{y}_{\mathrm{exp}}^{\text{rejected}}$, respectively. These ranked pairs form the basis for the preference optimization objective described in the next section.
Metric LLaMA3.1-8B LLaMA3.2-3B
ECQA (#10882) Acc. 71.11 65.85
CC-Cos Worst 07.70 01.30
Mean 08.18 01.65
Best 08.67 01.99
CC-Sp Worst 05.07 09.75
Mean 18.47 22.28
Best 31.76 34.64
ARC-E (#5197) Acc. 87.01 81.90
CC-Cos Worst 13.31 05.32
Mean 14.16 05.86
Best 15.02 06.41
CC-Sp Worst -01.06 07.26
Mean 12.44 19.68
Best 25.77 31.98
ARC-C (#2590) Acc. 77.57 68.70
CC-Cos Worst 15.09 07.67
Mean 16.07 08.29
Best 17.08 08.93
CC-Sp Worst -01.82 05.94
Mean 11.48 18.44
Best 24.73 30.89
Coda-H (#2776) Acc. 83.39 75.48
CC-Cos Worst 12.87 07.49
Mean 13.68 08.06
Best 14.52 08.64
CC-Sp Worst 07.11 06.82
Mean 19.97 19.39
Best 32.66 31.95
Table 1: Evaluation results across tasks. Acc. = Accuracy; Cos = CC-LIME-Cos; Sp = CC-LIME-Sp; (x100).

### v3

# Aligning What LLMs Do and Say: Towards Self-Consistent Explanations

@@ Abstract
Large language models (LLMs) seem to offer an easy path to interpretability: just ask them to explain their answers. Yet the features driving an answer often differ from those emphasized in its explanation, meaning post-hoc rationales can misrepresent what actually shaped the model’s output. We quantify this gap by comparing the feature-importance distributions of answers and their explanations. Prior analyses reveal such discrepancies, but large-scale study has been limited by the high computational cost of attribution methods. To address this, we introduce the Post-hoc Self-Consistency Bank (PSCB), a large-scale benchmark linking model decisions with diverse explanations and attribution vectors across datasets, methods, and model families. Using PSCB, we find that Spearman rank correlation provides a more reliable signal of alignment than cosine similarity. Building on this insight, we apply Direct Preference Optimization (DPO) to attribution-based preference data, improving alignment without degrading task accuracy, and show that standard supervised fine-tuning on the same data fails to achieve comparable gains. These improvements generalize robustly across domains, paving the way toward scalable and faithful alignment between LLM decisions and their natural language explanations. [^1] Code and data are available at https://github.com/saharad1/ConstLLM.
Aligning What LLMs Do and Say: Towards Self-Consistent Explanations
Sahar Admoni1, Ofra Amir1, Assaf Hallak2, Yftah Ziser2,3 1Technion – Israel Institute of Technology  2Nvidia Research  3University of Groningen saharad@campus.technion.ac.il, oamir@technion.ac.il, {ahallak,yziser}@nvidia.com
@@ 1 Introduction
As LLMs are increasingly embedded in user-facing and decision-support systems, their outputs and accompanying explanations are often trusted by end users even when independent verification is impractical Sun et al. (2024); Liu et al. (2023); Doshi-Velez and Kim (2017). To foster such trust, LLMs are frequently prompted to produce natural language explanations Madsen et al. (2024). Yet these post-hoc explanations often draw on different input features than those that determined the model’s answer Turpin et al. (2023); Randl et al. (2024); Hase et al. (2020), revealing a gap we term self-consistency. While faithfulness broadly concerns whether explanations reflect the model’s decision-making process Jacovi and Goldberg (2020); Wiegreffe et al. (2020), prior work suggests that many proposed faithfulness metrics in fact capture forms of self-consistency, agreement between the factors influencing a model’s prediction and those invoked in its explanation Parcalabescu and Frank (2024). We operationalize this notion through feature-attribution alignment, measuring whether the explanation relies on the same input evidence that influenced the model’s answer. Our work focuses on quantifying and improving this property for post-hoc explanations generated after the answer (see Figure 1).
Prior works apply counterfactual interventions to input features using various heuristics to estimate their influence on model decisions Wiegreffe et al. (2020); Turpin et al. (2023); Lanham et al. (2023); Atanasova et al. (2023). However, applying such interventions to LLMs is computationally expensive. Parcalabescu and Frank (2024) recently proposed a more rigorous approach based on feature importance, where self-consistency is defined as the similarity between the importance assigned to the answer and to its explanation. Yet, computing feature importance itself is resource-intensive, and their evaluation was limited to only 100 test examples, constraining the conclusions that can be drawn. To the best of our knowledge, no prior work has demonstrated how to improve this alignment, as even measuring it remains highly expensive.
We address this gap with the following contributions: (1) Post-hoc Self-Consistency Bank (PSCB): a large-scale benchmark linking over 85,000 decisions with 428,000 explanations and attribution vectors (LIME and LIG) across four QA datasets and two LLMs, enabling systematic evaluation of explanation–decision alignment. (2) Empirical analysis: The first large-scale study of attribution-based self-consistency, showing it is largely orthogonal to correctness and that Spearman rank correlation provides a more discriminative signal than cosine similarity, effectively separating high- and low-quality explanations. (3) Preference-based optimization: Attribution-based preference data from PSCB is used to fine-tune LLMs with DPO, yielding substantial in-domain gains and robust cross-domain generalization without degrading accuracy. (4) Multidimensionality: Improvements transfer across domains but not across attribution paradigms, revealing that different methods capture fundamentally distinct notions of input relevance, with direct implications for how the community evaluates attribution-based explanations.
@@
Figure 2: Overview of the PSCB pipeline. 1 Decision: Given a multiple-choice QA instance, the LLM generates an answer together with attribution scores over the input. 2 Explanation: The model is then prompted to produce $K$ diverse explanations (via temperature sampling) conditioned on the question and answer, with attribution scores computed for each explanation. 3 Alignment: Self-consistency scores are obtained by comparing attribution vectors from the decision and explanations. 4 Optimization: Finally, attribution-based preference pairs are used to fine-tune the model DPO, yielding more self-consistent explanations.
@@ 2 Background and Related Work
Feature attribution methods estimate how input components (e.g., tokens) or internal elements (e.g., attention heads) influence model predictions Zhao et al. (2023). Attention-based methods are simple but often unreliable (Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019), whereas gradient- and perturbation-based methods provide stronger signals, including Integrated Gradients Sundararajan et al. (2017), LRP Bach et al. (2015); Montavon et al. (2019), SHAP Lundberg and Lee (2017), and LIME Ribeiro et al. (2016). SHAP is rigorous but slow; LIME trades some precision for efficiency.
Despite advances in robustness and faithfulness Parcalabescu and Frank (2024); Atanasova et al. (2023), scalable, model-agnostic attribution remains challenging. LLMs often produce fluent yet unfaithful explanations (Narang et al., 2020; Turpin et al., 2023; Madsen et al., 2024), motivating self-consistency, alignment between features driving predictions and explanations. Perturbation-based tests evaluate robustness, while attribution-based methods compare importance vectors directly. Prior work such as CC-SHAP (Parcalabescu and Frank, 2024) handles only $\sim$100 samples; our Post-hoc Self-Consistency Bank (PSCB) scales this to tens of thousands via a LIME-based variant (CC-LIME). We also explore Layer-Integrated Gradients, a gradient-based method with cost between LIME and SHAP. Complementary efforts enhance explanation fidelity via probability-aware metrics (Siegel et al., 2024), context-faithful prompting (Zhou et al., 2023), or cross-example semantic coherence (Chen et al., 2024). Other recent work targets related but distinct objectives: PEX (Zhao and Iii, 2025) evaluates whether an explanation linguistically supports a label by comparing label-explanation log-odds, and CCT (Siegel et al., 2024) computes population-level correlations between counterfactual interventions and explanation mentions. Both operate at the level of probabilistic or semantic consistency rather than input-level attribution alignment. Because these methods do not access or compare token-level attribution vectors, they measure a complementary dimension of explanation quality; a model can score well on probabilistic consistency while still exhibiting misaligned feature attributions between its decision and explanation. Our work focuses on attributional self-consistency within each decision, directly comparing the input evidence used for the answer with that invoked in the explanation, and provides the first scalable framework for both measuring and improving this property.
3 Post-hoc Self-Consistency Bank (PSCB)
We present the Post-hoc Self-Consistency Bank (PSCB), a collection of attribution-augmented QA datasets for evaluating decision–explanation alignment. Each dataset is defined by three components: a base QA dataset, a target LLM, and an alignment metric. We describe our construction choices and report key statistics and findings. The pipeline (Figure 2) is agnostic to the choice of the attribution method and alignment metric.
@@ 3.1 Sequence Feature Attribution
We compute token-level attribution scores to analyze how input tokens influence a model’s generation. Given an input $\mathbf{x}=[x_{1},\dots,x_{m}]$ and an output $\mathbf{y}=[y_{1},\dots,y_{n}]$, the model generates $\mathbf{y}$ autoregressively with logits $\ell_{t}\in\mathbb{R}^{V}$ at each step $t$. We define the sequence log-probability as $\mathrm{SLP}(\mathbf{x},\mathbf{y})=\sum_{t=1}^{n}\log P\!\left(y_{t}\mid\mathbf{x},\mathbf{y}_{<t}\right)$, where $P\!\left(y_{t}\mid\mathbf{x},\mathbf{y}_{<t}\right)=\frac{\exp(\ell_{t}[y_{t}])}{\sum_{v\in V}\exp(\ell_{t}[v])}$, and $\ell_{t}$ denotes the model logits at step $t$ conditioned on $(\mathbf{x},\mathbf{y}_{<t})$. This scalar serves as the attribution target, indicating how much each token in $\mathbf{x}$ contributes to generating $\mathbf{y}$. For each output $\mathbf{y}$, we compute an attribution vector $\boldsymbol{\phi}^{(\mathbf{y})}=[\phi_{1}^{(\mathbf{y})},\dots,\phi_{m}^{(\mathbf{y})}]$, where $\phi_{i}^{(\mathbf{y})}$ quantifies the contribution of $x_{i}$. The attribution method, gradient-based, perturbation-based, or sampling-based, determines how $\boldsymbol{\phi}^{(\mathbf{y})}$ is computed; our framework is method-agnostic (Section 3.4). To reduce noise, we exclude a set of skip tokens $\mathcal{S}\subseteq V$ (e.g., punctuation, formatting markers), ensuring attribution focuses on semantically meaningful inputs.
@@ 3.2 Measuring Self-Consistency
Building on the attribution vectors defined above, we evaluate whether an explanation reflects the reasoning behind a decision. For an input $\mathbf{x}$ with decision $\mathbf{y}_{\mathrm{dec}}$ and explanation $\mathbf{y}_{\mathrm{exp}}$, let $\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})}$ and $\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}})}$ denote their respective attributions. Self-consistency is measured by an alignment function
$\alpha\!\left(\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})},\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}})}\right):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R},$
where the choice of $\alpha$ determines the aspect of the alignment to be evaluated (e.g. magnitude, rank, or overlap). Specific metrics are detailed in the following subsections.
@@ 3.3 Constructing Attribution-Based Preference Data
Building on attribution vectors (Section 3.1) and alignment measures (Section 3.2), we construct a dataset that pairs model outputs with their attributions, allowing systematic comparison between a model’s decision rationale and its post-hoc justifications. Each instance is represented as a quintuple
$\big(\mathbf{x},\;\mathbf{y}_{\mathrm{dec}},\;\{\mathbf{y}_{\mathrm{exp}}^{(i)}\}_{i=1}^{k},\;\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{dec}})},\;\{\boldsymbol{\phi}^{(\mathbf{y}_{\mathrm{exp}}^{(i)})}\}_{i=1}^{k}\big),$
where $\mathbf{x}$ is the input, $\mathbf{y}_{\mathrm{dec}}$ the predicted answer, and $\{\mathbf{y}_{\mathrm{exp}}^{(i)}\}_{i=1}^{k}$ are diverse explanations sampled in a zero-shot setting by conditioning on $(\mathbf{x},\mathbf{y}_{\mathrm{dec}})$. The corresponding attribution vectors $\boldsymbol{\phi}$ are computed as in Section 3.1, with non-semantic tokens excluded. We then score each explanation by its alignment $\alpha$ with the decision (Section 3.2), and retain the highest- and lowest-scoring cases as $\mathbf{y}_{\mathrm{exp}}^{\text{chosen}}$ and $\mathbf{y}_{\mathrm{exp}}^{\text{rejected}}$. These preference pairs form the core of the dataset used for optimization (Section 4).